OpenStudy (khally92):

A radio active substance has a half life of 20-days how much time is required so that only 1/32 of the original amount remains? Find the rate of decay at this time..

OpenStudy (anonymous):

100 days because 1/32=3.125% 100/2=50 1st half life 50/2=25 2nd half life 25/2=12.5 3rd half life 12.5/2=6.25 4th half life 6.25/2=3.125% 5th half life and since each half life is 20 days 20*5=100 days

OpenStudy (anonymous):

In other words, \[\frac{1}{32}=\left(\frac{1}{2}\right)^5\] 5 half-lives = 5×20days.

OpenStudy (khally92):

yeah better..

OpenStudy (khally92):

what about the rate of decay/

OpenStudy (anonymous):

That you need to solve for using logarithms. How are you with taking logs of stuff?

OpenStudy (khally92):

am good

OpenStudy (anonymous):

Sweet. Here's how it goes: Half-life is 20days, so we'll start with t=20. \[A(t)=A_0e^{kt}.\] (General exponential growth/decay model) Half-life means that A(20)=0.5A_0 -> \[\frac{1}{2}=e^{20k}.\] Now take natural log of both sides.

OpenStudy (anonymous):

After solving for k, you can put it back in the equation for \[\frac{1}{32}=e^{kt}\] and solve for t to verify that it is 100 days.

OpenStudy (anonymous):

Clear so far, @khally92 ?

OpenStudy (khally92):

yeah sure thanks...

OpenStudy (anonymous):

OK, you should get something around -0.035 for the decay rate.